OpenStudy (anonymous):

can someone walk me through this to me please? polar coordinates confuse me. i know how to do double integrals, but not how to determine bounds within a polar and what to sub where. Using polar coordinates, evaluate the integral ∫∫sin(x^2+y^2)dA where R is the region 16≤x^2+y^2≤81.

zepdrix (zepdrix):

Hey there :) So recall that x^2+y^2 represents a circle, yah?

zepdrix (zepdrix):

So the region that we're integrating over is the space between two circles.

zepdrix (zepdrix):

We want to find the area between these two circles. Understand how I was able to come up with those radius values?

zepdrix (zepdrix):

To convert to polar we use these:\[\Large\rm x=r \cos \theta\]\[\Large\rm y=r \sin \theta\]Recall that when you change from Cartesian to Polar, your differentials will pick up a factor of r,\[\Large\rm dA=\quad dy~dx=\quad r~dr~d\theta\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{6}^{81}\sin(r^2)drd\]??

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

One of our boundaries is this:\[\Large\rm 16\le x^2+y^2\]\[\Large\rm 4^2\le x^2+y^2\]Recall your circle equation:\[\Large\rm x^2+y^2=r^2\]We need to see that value squared so we can get a proper boundary for \(\Large\rm r\).

zepdrix (zepdrix):

So we want our radius to extend from r=4 out to r=9.

zepdrix (zepdrix):

And, uh oh... you forgot your factor of r when you converted your differentials!!! :O

OpenStudy (anonymous):

ok, so it's \[\int\limits_{4}^{9}\], not 16, 81